# From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks

### v1

@@ 2.1 From Frege to Fodor
Two further consequences of the Chomskyan revolution deserve mention before concluding this historical overview. First, the rise of this new explanatory paradigm brought with it a notable resurgence of nativism, as Chomsky and others argued that the linguistic data to which children are exposed is insufficient for recovering the kind of fine-grained syntactic structures necessary for explaining human linguistic competences (Laurence and Margolis,, 2001). As we will briefly discuss in 5.2, the rise of DNNs and LLMs may make novel contributions to this long-standing philosophical issue (Buckner,, 2018, 2019; Buckner, 2023b, ). And second, though we have framed the discussion in terms of the meaning of sentences throughout this section, this period saw a distinct shift in focus from language to thought. This widely discussed ‘cognitive turn’ (Miller,, 2003) overthrew behaviorist strictures on psychological explanation (Fodor, 1968b, ) and transposed many of the debates central to the linguistic turn into questions about the structure of cognition itself (Burge,, 2007). Compositionality was no exception. Jerry Fodor influentially argued that the compositional structure of natural languages is a reflection of the compositional structure of thought, reintroducing the language of thought hypothesis to the philosophical mainstream (Fodor,, 1975; Pelletier and Roques,, 2017). On this view, human cognition is compositional from the ground up—it is a feature of the mind’s basic operating system, not just an artifact of one specific domain (language).
@@
Across this historical evolution, compositionality arrives in the position familiar to cognitive scientists today: an empirically motivated, non-negotiable property of both language and thought and the best explanation of the creativity, systematicity and productivity of human cognition. Throughout his career, Fodor would wield compositionality like a sledgehammer to smash rival theories of meaning and defend the language of thought hypothesis. In fact, he essentially runs the contrapositive of Ryle’s argument above—because the compositionality principle is non-negotiable and the use theory of meaning does not ensure it, the use theory of meaning must be false (Fodor and Lepore,, 2001; Fodor,, 2003). Over the years, he employed the same logic to attack conceptual role semantics (Fodor and LePore,, 1993), prototype theory (Fodor and Lepore,, 1996), and most importantly for our present purposes, to neural network models of cognition (Fodor and Pylyshyn,, 1988).
@@ 2.2 The Rise of Connectionism
In the 80s and 90s, a series of key advances in neural networks, such as the invention of Hopfield networks (Hopfield,, 1982) and the backpropagation algorithm (Rumelhart et al., 1986a, ), led many to abandon this classical, Fodorian picture and turn to an alternative view of the mind (McClelland et al., 1986b, ; Rumelhart et al., 1986b, ). The connectionist movement was initially motivated by considerations other than compositionality, such as the basic computational structure of biological brains. Proponents sought to demonstrate that sophisticated cognitive behaviors could emerge in networks of simple interconnected units (or “neurons”). Though these units abstract away from a great deal of what is known about biological neurons, they implement a similar kind of computation: each unit performs a weighted sum of its inputs and applies a threshold or nonlinearity, much as biological neurons integrate dendritic inputs and spike at a certain threshold (Hodgkin and Huxley,, 1952; McCulloch and Pitts,, 1943; O’Reilly et al.,, 2012). Early work demonstrated how these networks could learn by incrementally updating their weights (or “synapses”) to improve overall performance on a given task (Rumelhart et al., 1986b, ).
@@ 2.3 The Compositionality Challenge
This first horn of Fodor and Pylyshyn’s argument draws on two key cognitive properties that are taken to require compositionality: productivity and systematicity. As we have seen, classical symbolic accounts provide a natural mechanism for productively generating an unbounded set of novel thoughts through the recombination of a finite pool of primitive elements. Connectionist architectures, on the other hand, do not seem to build in the kind of combinatorial constituent structure capable of supporting this form of unbounded productivity. They also struggle to capture the systematicity of cognition—the apparently intrinsic connection between the ability to comprehend and generate specific sets of thoughts/sentences. To use the classic example, it seems that anyone capable of understanding the sentence “John loves Mary” should be capable of understanding the sentence “Mary loves John.” The classical account again provides a natural explanation—the meaning of the second sentence is a function of the very same components and mode of combination as the first. Therefore, any compositional learner capable of understanding the first will be able to redeploy this knowledge in understanding the second. [^11] This systematicity challenge differs from the challenge provided by productivity. For one thing, it does not rely on idealizing to an infinite cognitive capacity, and so may seem to yield the same conclusion from a weaker premise (a key motivation for introducing the argument in the first place—see Fodor,, 1987, p. 147-154). For important qualifications, however, see Szabó, (2022). Fodor and Pylyshyn argue that in a connectionist network without constituent structure, on the other hand, each of these individual acts of understanding would be an independent capacity. There would therefore be no inherent reason why they should hang together—a network trained on one could easily misunderstand the other. For both productivity and systematicity, Fodor and Pylyshyn’s conclusions are the same—without underlying compositional mechanisms, connectionist networks will fail to replicate these core features of human language and cognition.
@@
The second horn of Fodor and Pylyshyn’s dilemma draws on the possibility of implementing a classical symbolic system in an artificial neural network. Even if a connectionist network did manage to exhibit systematic and productive behaviors, they argue, this would have to be because it had merely implemented a classical symbolic system, undermining any claim to revolutionary implications or novel insight into the nature of the mind (for further discussion, see section 5.3). Thus, the connectionist is left with no way out: either their models are empirically inadequate, or they are mere implementations, which may prove informative for theorists interested in how symbolic systems are realized in the human brain (McLaughlin,, 1993), but will not contribute to higher-level cognitive theory.
As should be clear, this argument is a species of the genus described in section 2.1—compositionality is treated as an empirically motivated psychological construct, necessary for the explanation of certain key human behaviors.
In the context of connectionism, however, the compositionality challenge ignited a heated debate and drew a wide array of rebuttals and clarifications (van Gelder,, 1990; Matthews,, 1994; Ramsey et al.,, 1990; Horgan and Tienson,, 1991). Some connectionist researchers—most notably Paul Smolensky—tried to meet this ‘mere implementation’ objection head on by building neural networks with internal compositional structure that did not merely recapitulate classical architectures (Smolensky,, 1990, 1991). This led to further debate about just what kind of explanation compositionality offered and what internal compositional structure was rightly considered ‘non-negotiable’ (MacDonald and MacDonald,, 1991; van Gelder,, 1990). Robert Hadley, (1997) observed that the demand that systematicity be a matter of nomic necessity, an inherent or intrinsic element of the architecture, is too strong, since even a classical symbolic architecture could be programmed to exhibit failures of systematicity. And David Chalmers, (1993) pointed out that Fodor and Pylyshyn’s argument proves too much and so cannot be correct, as formulated. As all parties to the debate acknowledge, connectionist networks may implement classical symbolic architectures. However, Chalmers noted, Fodor and Pylyshyn’s argument for the first horn of the dilemma hinges on basic features of neural networks that are present even in such connectionist implementations of classical architectures. Yet nowhere do they specify an ‘escape clause’ by which these implementations are exempt from the first horn of the argument (Chalmers,, 1993). Thus, if their initial argument were correct, it would prove that even classical symbolic architectures (implemented in neural networks) lack productivity and systematicity, a conclusion that is clearly untenable and self-defeating.
Our primary goal in this paper—and the sole focus of the review in section 4—is to survey recent findings suggesting that contemporary DNNs have plausibly surmounted the first horn of this dilemma. In 5.3, we also briefly evaluate whether this means that they fall on the second horn (for further discussion, see McGrath et al., 2023a, ; McGrath and Russin,, 2024).
@@
3 Neural Networks in the Age of Deep Learning
@@
The primary factor that differentiates modern neural networks from their connectionist predecessors is their scale: modern networks are deeper (i.e. more layers), have many orders of magnitude more learnable parameters (weights), and are trained on many orders of magnitude more data. This is possible due to a combination of advances in hardware (e.g., GPUs) and the availability of large amounts of data from the internet (e.g., Wikipedia). These are merely quantitative differences, but they have led to qualitative increases in their capabilities (Brown et al.,, 2020; Caballero et al.,, 2023; Kaplan et al.,, 2020; Wei et al., 2022b, ).
While scale has perhaps been the single most important factor, there have also been advances in areas such as network architecture, where the most notable development has been the introduction of the transformer (Vaswani et al.,, 2017). In this section, we will briefly review transformers. [^12] Readers already familiar with transformers can skip directly to section 4. Those interested in learning even more about transformers may want to read the original paper in which they were introduced (Vaswani et al.,, 2017), as well as helpful introductory texts (e.g., Millière and Buckner, 2024a, ), and some of the many helpful blog posts available online (e.g., Alammar,, 2018).
Invented in 2017, the transformer has become the dominant architecture in natural language processing models (e.g., in the LLMs powering chatGPT) and has also become common in vision models (Dosovitskiy et al.,, 2021; Dehghani et al.,, 2023), multimodal models (Xu et al.,, 2023), and even reinforcement learning models (Li et al.,, 2023). Transformers take a sequence of vectors as inputs and produce a sequence of vectors as outputs. In the case of language models, each input vector will be an embedding of an individual word (or to be more precise, a sub-word token, which can be a whole word or a part of a word; Sennrich et al.,, 2016; Radford et al.,, 2019). One of the main advantages of the transformer architecture is that it processes input sequences (e.g., sentences, paragraphs, or documents of text) in parallel, avoiding the computational cost of recurrent neural networks (RNNs), which need to compute intermediate states serially across the sequence (see Figure 1A). RNNs maintain and update a single hidden state as the input sequence is processed serially, effectively requiring them to continuously compress the entire past (i.e., anything from prior inputs that could be relevant to future outputs) into this single hidden state. Transformers have no such constraint, as encodings at each position in the sequence can be updated by information in arbitrary other positions via attention (see Figure 1B).
Figure 1: Transformer basics. A) Comparison with recurrent neural networks (RNNs) in the language modeling setting (predicting the next word). A fundamental difference between RNNs and transformers is that RNNs must compress the entire past into a single hidden vector. To see this, notice how the only way information from the word ‘Attention’ can get to the position of the word ‘you’ is via the hidden states at the intermediary positions (highlighted with red arrows). In the transformer, information from the word ‘Attention’ can get to the position of the word ‘you’ directly at any layer via the attention mechanism (single red arrow). Attention weights are depicted by the thickness/darkness of the lines (which have been chosen arbitrarily for illustration purposes). B) Attention mechanism. At each position in the sequence, queries (blue), keys (green) and values (purple) are determined by multiplying inputs by learnable weight matrices ($W_{q}$ , $W_{k}$ , $W_{v}$). For simplicity, key and value are not shown at position 4, and queries are not shown at other positions. Queries (blue) are compared to keys (green), and the result determines the attention weight. A good match means the value vector (purple) at that key’s position will receive a large weight (shown by darker lines).
The attention mechanism (Bahdanau et al.,, 2014; Vaswani et al.,, 2017) is so named because it allows representations at a given position in the sequence to selectively filter (“attend to”) the information it receives from other positions (like how humans selectively filter information in the environment in order to accomplish a goal; Lindsay,, 2020). In self-attention, a layer of encodings of a sequence can attend to the previous layer of encodings of the same sequence. This is opposed to cross-attention, where a decoder attends to a given sequence while producing a new sequence (e.g., in another language during a translation task). Each layer of the transformer has a number of heads, which process the input vector at each position in the sequence, multiplying it by three separate learnable matrices to form a query, a key, and a value vector. Intuitively, the query vectors can be thought of as soliciting information that is needed at a given position (e.g., what is this verb’s direct object?), the key vectors as advertising what information is available at a given position (e.g., that this noun might be a direct object), and the value vectors as representing the content that will actually get transmitted if there is a good match between key and query (e.g., that this direct object is “John”). Thus, the keys and the queries fully determine how information will be transmitted across positions in the sequence according to how well they match, as measured by their dot product:
$\text{Attention}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d}})V$ (1)
Here, the matrix multiplication $QK^{T}$ computes a matrix where each element is the dot product of a particular query-key pair (normalized by the square root of the dimension of the vectors, $d$), and the softmax ensures that each row of the matrix will be a probability distribution, which determines the weights that multiply the value vectors in $V$. The final output at a given position of one attention head, then, is an attention-weighted average of the value vectors from all positions, where the attention weights are determined by how good the match is between that position’s query and another position’s key. [^13] By default, the “residual stream” (Elhage et al.,, 2021) of the transformer does not mix information across positions; for information to get from position i to position j, the key vector at position i must have a large dot product with the query vector at position j. This will give a large value for the input to the softmax at entry (i,j) which will mean a higher proportion of the final output at position j will be constituted by the value vector from position i. In effect, the attention mechanism allows the network to combine and integrate information across the sequence in arbitrarily flexible yet selective ways.
Importantly, the designer of a transformer doesn’t actually specify in advance how the attention should operate (i.e., what kinds of information will go into the queries, keys, or values for a given head)—these will be resolved in the course of the optimization process, where the network will discover the best ways of using the learnable parameters in the matrices of the attention mechanism to succeed on its task (i.e., to reduce its loss). Some recent work has attempted to understand what kinds of representations and attentional strategies these networks end up discovering. For example, some attention heads in language models seem to be sensitive to syntactic structures (Manning et al.,, 2020). Other work has found “induction heads” (Olsson et al.,, 2022) that search the context for previous occurrences of the current token and encode what came after it, thus allowing the network to “reason by induction” to predict that the same pattern will repeat again. Rapid progress has been made in methods for discovering the circuits underlying a given capacity in trained LLMs (“mechanistic interpretability”; Conmy et al.,, 2023; Elhage et al.,, 2021; Merullo et al.,, 2024; Olsson et al.,, 2022; Wu et al.,, 2023), but it is an important open question whether the mechanisms learned by LLMs in the course of training will all afford these kinds of succinct algorithmic descriptions (Arrieta et al.,, 2019; Crook et al.,, 2023; Lipton,, 2017).
Thus, the fundamental spirit of connectionism is alive and well in the deep learning revolution: practitioners don’t build into their networks the representations and processes that they will use to solve a given task, instead allowing useful representations to emerge in their distributed patterns of activity learned through trial and error on data. In some ways, modern deep learning researchers have taken this idea to new extremes. Whereas older work in machine learning used highly engineered features to make the learning problem easier for a downstream learning system such as a shallow neural network, modern practitioners skip the feature-engineering step entirely and train deep networks directly on raw images or text, allowing the early layers to discover whatever features will be most useful for downstream layers through “end-to-end” learning (Goodfellow et al.,, 2016). This “bitter lesson” (Sutton,, 2019)—that in high-dimensional domains like vision or language, deep neural networks will usually find better solutions than the ones that even the cleverest engineers could come up with—stands in direct conflict with the spirit of Chomskyan and Fodorian nativism, which emphasizes that certain properties (e.g., compositionality) must be necessary consequences of the specific kinds of representations that are built into the architecture (Fodor and Pylyshyn,, 1988; Fodor,, 1975).
4 Review
@@
Although these recent developments in DNNs have led to remarkable performance gains across many domains, it is unclear a priori whether any of the specific advances (e.g., increasing scale, the transformer architecture) should have a significant impact on their underlying compositionality. The features of neural networks that led theorists like Fodor and Pylyshyn, (1988) to conclude that they were not compositional seem prima facie to be largely unchanged in modern deep networks: transformers are still extremely unbiased learners with no explicit rule-like representations or built-in sensitivity to constituent structure. [^14] While most agree that the transformer is a very unbiased architecture (e.g., it is in some ways even more unbiased than the convolutional neural network; Dosovitskiy et al.,, 2021), some have argued that the attention mechanism might play a special role in the compositional behaviors observed in transformers (Smolensky et al., 2022a, ). As we have seen, accounting for compositionality is a pressing theoretical challenge—if neural networks are to provide illuminating accounts of human cognitive capacities, they must be able to model the creative, productive, and systematic generalization behaviors exhibited by humans.
@@
Our central question in this review will be an empirical one: can contemporary neural networks replicate the behavioral signatures of compositionality? We will argue that although many current neural networks (including ones leveraging modern architectures such as the transformer) still seem incapable of capturing these behavioral signatures, a few novel approaches have shown that it is possible to endow neural networks with the properties that are required to produce them. These results cast doubt on the empirical predictions made by classical theorists that neural networks would never be capable of generalizing systematically or productively, and suggest that DNNs may meet the challenge posed in the first horn of Fodor and Pylyshyn’s (1988) dilemma.
@@ 4.1 Operationalizing Compositionality
One might think that in order to determine whether a given learner is really compositional, its internal representations and/or processes need to be scrutinized in addition to its behavior. For example, some work has attempted to measure the extent to which a model’s internal representations capture ground-truth compositional or constituent structure (Andreas,, 2019; Lepori et al.,, 2023; Lewis et al.,, 2023; McCoy et al.,, 2019; Russin et al., 2021a, ; Soulos et al.,, 2020; Todd et al.,, 2024). We think this work is interesting and briefly discuss it in the final section (5.3). However, we see this interpretive work as being relevant not to the first horn of the compositionality challenge, but to the second horn—the question of whether a system is producing compositional behaviors by ‘merely implementing’ a classical architecture. As we discussed in section 2.3, the first horn of the dilemma, in our view, is an empirical question that can be answered with reference to behavioral evidence, while the second horn of the dilemma requires both empirical investigation into the inner workings of a given neural network and philosophical work to determine what counts as a ‘classical architecture’ (Garson,, 1994), whether a given network is implementing one, and what should count as a satisfactory explanation of systematic and productive behavior.
@@
Figure 2: Are neural networks stimulus-response learners, or do they learn to solve problems with compositional algorithms? Two extreme possibilities are depicted for natural language (top) and simple arithmetic (bottom). In either case, a stimulus-response learner (left) would simply memorize a lookup table where entries corresponded to entire inputs/outputs, ignoring any compositional structure. A compositional learner (right) would solve the same problems by composing familiar elements according to known rules, allowing generalization. Figure reproduced from Russin et al., 2021a .
@@
In a seminal study on compositionality in modern DNNs, Lake and Baroni, (2018) leveraged these principles to develop the SCAN task, a sequence-to-sequence task where the model takes instructions as inputs (e.g., “walk left”), and must return the corresponding sequence of actions (e.g., TURN_LEFT WALK). The instructions were generated from a simple, finite phrase-structure grammar (about 20,000 total possible instruction sequences), and corresponding action sequences were generated by a deterministic interpretation function. To test compositional generalization, the researchers created a systematic difference between the train and test sets: one of the four primitive verbs (“jump”) was left out of the training set except in its simplest form (“jump” → JUMP). [^17] The dataset also included other generalization tests, such as a split designed to test generalization to longer sequences (Lake and Baroni,, 2018). Here we focus on the ‘add-jump’ split, which was the main compositional generalization test on which standard networks were shown to fail. This created a scenario like the one discussed above, where a model could in principle infer the underlying grammar from all of the complicated constructions containing the other primitive verbs (e.g., “walk twice,” “look twice,” etc.), and combine this with knowledge of the action associated with “jump” (JUMP) to generalize to the analogous constructions containing “jump” (e.g., “jump twice”).
@@
The results showed that even though standard modern DNNs performed nearly perfectly when the train and test sets were randomly split (i.e., when the train and test data were randomly sampled from all possible instructions), these same networks performed quite poorly on this compositional generalization test (about 1.2% accuracy reported for the RNNs tested in the original study). This was true for all of the models they tested, and follow-up work showed that other standard architectures such as convolutional neural networks (Dessì and Baroni,, 2019) and transformers (Keysers et al.,, 2020) exhibit similar reductions in performance on the compositional test (though less extreme in the case of the convolutional neural network). Analogous experiments in humans using a modified version of the same task have shown that humans perform quite well (Lake et al.,, 2019; Lake and Baroni,, 2023). A number of subsequent machine learning studies followed up on this work, creating similar tests of compositional generalization and showing that these standard DNNs fail to generalize compositionally in domains such as semantic parsing (Kim and Linzen,, 2020), string manipulation with a probabilistic context-free grammar (Hupkes et al.,, 2020), mathematics (Saxton et al.,, 2019), natural-language translation (Dankers et al.,, 2022), and grounded language understanding (Ruis et al.,, 2020), among others.
@@
Figure 3: Two different kinds of generalization in one dimension (top) and in higher dimensions, as in the SCAN task (bottom). When train data and test data come from the same domain, generalization to the test data only requires interpolation (top left). When test data is sampled from a different domain, generalization requires extrapolation (top right). In the bottom row, each shape represents the distribution of all possible instruction sequences in the SCAN grammar (Lake and Baroni,, 2018), which can be split into train and test in different ways. When train and test data are sampled from the same distribution, i.e., independent and identically distributed (i.i.d.), neural networks generalize well (bottom left). However, in compositional generalization tests there are systematic differences between train and test data, thus requiring out-of-domain (o.o.d.) generalization. For example, in the compositional generalization test in the SCAN task (Lake and Baroni,, 2018), the instructions containing “jump” are systematically left out of the training set (bottom right). In reality these samples would be difficult to visualize in two dimensions, but for the sake of illustration, all such instructions are depicted here as falling in the upper left quadrant of the distribution to emphasize the extrapolation required in the o.o.d. case. Additionally, each quadrant is depicted as sharing an underlying symmetry with the other quadrants, corresponding to the symmetry with the set of instructions containing the other three primitive verbs. Figure adapted from (Russin et al., 2020a, ).
@@
Another (perhaps compatible) way of understanding these results is to take a statistical learning theory perspective (see Figure 3). Neural networks are universal function approximators (Cybenkot,, 1989; Hornik,, 1991) that assume that the train data and test data are sampled from the same distribution, i.e., that the train and test data will be independent and identically distributed (i.i.d.). Viewed through this lens, the generalization behaviors that are taken to be diagnostic of compositionality (Evans,, 1985; Fodor and Pylyshyn,, 1988) are out-of-domain (o.o.d.) generalizations (see Figure 3), and require extrapolation, rather than interpolation (Marcus,, 1998, 2018). [^18] Out-of-domain generalization is also sometimes referred to as “out-of-distribution” or “out-of-sample” generalization, and is required in any “non-stationary” environment where the data-generating distribution changes over time. Also, see Balestriero et al., (2021) for problems with this usage of the terms ”interpolation” and ”extrapolation” in high-dimensional settings. For an unbiased learner like a standard neural network, this kind of generalization should be unlikely or impossible.
@@
The extent to which a learner is expected to generalize beyond its training distribution is the extent to which it must make additional assumptions about the world. These assumptions, called ‘inductive biases’ in statistical learning parlance, can take many forms in neural networks, including architectural inductive biases (e.g., convolutional neural networks assume translation invariance/equivariance), extra regularization terms in the loss function (e.g., weight decay, a regularization term that encourages a network’s weights to be small), etc. Standard neural networks do not make very many of these assumptions and thus tend to struggle when even the most superficial distributional shifts occur between train and test (Lake et al.,, 2017). For example, standard neural networks used in deep reinforcement learning agents trained on Atari games fail to generalize when superficial or trivial features of the game are changed (Kansky et al.,, 2017). It is not surprising, then, that they also fail in compositional generalization settings where there can be extreme distributional shifts from train to test.

### v2

@@ 2.1 From Frege to Fodor
Two further consequences of the Chomskyan revolution deserve mention before concluding this historical overview. First, the rise of this new explanatory paradigm brought with it a notable resurgence of nativism, as Chomsky and others argued that the linguistic data to which children are exposed is insufficient for recovering the kind of fine-grained syntactic structures necessary for explaining human linguistic competences (Laurence and Margolis,, 2001). As we will briefly discuss in 4.2, the rise of DNNs and LLMs may make novel contributions to this long-standing philosophical issue (Buckner,, 2018, 2019; Buckner, 2023b, ). And second, though we have framed the discussion here in terms of the meaning of sentences, this period saw a distinct shift in focus from language to thought. This ‘cognitive turn’ transposed key debates about language into questions about the structure of cognition itself (Burge,, 2007). Compositionality was no exception. Jerry Fodor influentially argued that the compositional structure of natural languages is a reflection of the compositional structure of thought, reintroducing the language of thought hypothesis to the philosophical mainstream (Fodor,, 1975; Pelletier and Roques,, 2017). On this view, human cognition is compositional from the ground up—it is a feature of the mind’s basic operating system, not just an artifact of one specific domain (language).
@@
Across this historical evolution, compositionality arrives in the position familiar to cognitive scientists today: a core property of human cognition, motivated by its non-negotiable role in explaining the creativity, systematicity and productivity of language and thought. From this perspective, the inability to adequately account for compositionality is a fatal flaw, which Fodor wielded like a sledgehammer to smash rival theories of meaning and defend his language of thought hypothesis. Over the years, he would take this sledgehammer to the use theory of meaning (Fodor and Lepore,, 2001; Fodor,, 2003), conceptual role semantics (Fodor and LePore,, 1993), prototype theory (Fodor and Lepore,, 1996), and most importantly for our present purposes, neural network models of cognition (Fodor and Pylyshyn,, 1988).
@@ 2.2 The Rise of Connectionism
In the 80s and 90s, a series of key advances in neural networks, such as the invention of Hopfield networks (Hopfield,, 1982) and the backpropagation algorithm (Rumelhart et al., 1986a, ), led many to abandon the classical, Fodorian picture and turn to an alternative view of the mind (McClelland et al., 1986b, ; Rumelhart et al., 1986b, ). The connectionist movement was initially motivated by considerations other than compositionality, such as the basic computational structure of biological brains. Proponents sought to demonstrate that sophisticated cognitive behaviors could emerge from a network of simple interconnected units (or “neurons”). Though these units abstract away from a great deal of what is known about biological neurons, they implement a similar kind of computation: each unit performs a weighted sum of its inputs and applies a threshold or nonlinearity, much as biological neurons integrate dendritic inputs and spike at a certain threshold (Hodgkin and Huxley,, 1952; McCulloch and Pitts,, 1943; O’Reilly et al.,, 2012). Early work demonstrated how these networks could learn by incrementally updating their weights (or “synapses”) to improve overall performance on a given task (Rumelhart et al., 1986b, ).
@@ 2.3 The Compositionality Challenge
This first horn of Fodor and Pylyshyn’s argument draws on two key cognitive properties that are taken to require compositionality: productivity and systematicity. As we have seen, classical symbolic accounts provide a natural mechanism for productively generating an unbounded set of novel thoughts through the recombination of a finite pool of primitive elements. Connectionist architectures, on the other hand, do not seem to build in the kind of combinatorial constituent structure capable of supporting this form of unbounded productivity. They also struggle to capture the systematicity of cognition—the apparently intrinsic connection between the ability to comprehend and generate specific sets of thoughts/sentences. To use the classic example, it seems that anyone capable of understanding the sentence “John loves Mary” should be capable of understanding the sentence “Mary loves John.” The classical account again provides a natural explanation—the meaning of the second sentence is a function of the very same components and mode of combination as the first. Therefore, any compositional learner capable of understanding the first will be able to redeploy this knowledge in understanding the second. Fodor and Pylyshyn argue that in a connectionist network without constituent structure, on the other hand, each of these individual acts of understanding would be an independent capacity. There would therefore be no inherent reason why they should hang together—a network trained on one could easily misunderstand the other. For both productivity and systematicity, Fodor and Pylyshyn’s conclusions are the same—without underlying compositional mechanisms, connectionist networks will fail to replicate these core features of human language and cognition.
@@
The second horn of Fodor and Pylyshyn’s dilemma draws on the possibility of implementing a classical symbolic system in an artificial neural network. Even if a connectionist network did manage to exhibit systematic and productive behaviors, they argue, this would have to be because it had merely implemented a classical symbolic system, undermining any claim to revolutionary implications or novel insight into the nature of the mind (for further discussion, see section 4.3). Thus, the connectionist is left with no way out: either their models are empirically inadequate, or they are mere implementations, which may prove informative for theorists interested in how symbolic systems are realized in the human brain (McLaughlin,, 1993), but will not contribute to higher-level cognitive theory. As should be clear, this argument is a species of the genus described in section 2.1—compositionality is treated as an empirically motivated psychological construct, necessary for the explanation of certain key human behaviors.
The compositionality challenge ignited a heated debate and drew a wide array of rebuttals and clarifications (Chalmers,, 1993; Hadley,, 1997; Horgan and Tienson,, 1991; Matthews,, 1994; Ramsey et al.,, 1990). Some connectionist researchers—most notably Paul Smolensky—tried to meet the ‘mere implementation’ objection head on by building neural networks with internal compositional structure that did not merely recapitulate classical architectures (Smolensky,, 1990, 1991). This led to further debate about just what kind of explanation compositionality offered and what internal compositional structure was rightly considered ‘non-negotiable’ (MacDonald and MacDonald,, 1991; van Gelder,, 1990). Our primary goal in this paper—and the sole focus of the review in section 3—is to survey recent findings suggesting that contemporary DNNs have plausibly surmounted the first horn of this dilemma. In 4.3, we also briefly evaluate whether this means that they fall on the second horn (for further discussion, see McGrath et al., 2023a, ; McGrath and Russin,, 2024).
@@
3 In-Context Compositionality
@@
Many advances in neural networks have accompanied the deep learning revolution, including the introduction of the transformer architecture (Vaswani et al.,, 2017). However, the single most important factor that differentiates modern neural networks from their connectionist predecessors is their scale: modern networks are deeper (i.e. more layers), have many orders of magnitude more learnable parameters (weights), and are trained on many orders of magnitude more data. This is possible due to a combination of advances in hardware (e.g., GPUs) and the availability of large amounts of data from the internet (e.g., Wikipedia). These are merely quantitative differences, but they have led to qualitative increases in these systems’ capabilities (Brown et al.,, 2020; Caballero et al.,, 2023; Kaplan et al.,, 2020; Wei et al., 2022b, ).
@@
Although these developments have led to remarkable gains in performance across many domains, it is unclear a priori whether they should have a significant impact on DNNs’ underlying compositionality. The features of neural networks that led theorists like Fodor and Pylyshyn to conclude that they were not compositional seem prima facie to be largely unchanged in modern deep networks: DNNs are still extremely unbiased learners with no explicit rule-like representations or built-in sensitivity to constituent structure. [^9] While most agree that the transformer is a very unbiased architecture (e.g., it is in some ways even more unbiased than the convolutional neural network; Dosovitskiy et al.,, 2021), some have argued that the attention mechanism might play a special role in the compositional behaviors observed in transformers (Smolensky et al.,, 2022). As we have seen, accounting for compositionality is a pressing theoretical challenge—if neural networks are to provide illuminating accounts of human cognitive capacities, they must be able to model the creative, productive, and systematic generalization behaviors exhibited by humans.
@@
Our central question in this review will be an empirical one: can contemporary neural networks replicate the behavioral signatures of compositionality? We will argue that although many current neural networks (including ones leveraging modern architectures such as the transformer) still seem incapable of capturing these behavioral signatures, novel approaches have shown that it is possible to endow neural networks with the properties that are required to produce them. In particular, recent work has shown that metalearning neural networks, which learn how to learn new tasks given in context, can reproduce key compositional generalization behaviors. These results cast doubt on the empirical predictions made by classical theorists that neural networks would never be capable of generalizing systematically or productively, and suggest that DNNs may meet the challenge posed in the first horn of Fodor and Pylyshyn’s dilemma.
@@ 3.1 Operationalizing Compositionality
Figure 1: Are neural networks stimulus-response learners, or do they learn to solve problems with compositional algorithms? Two extreme possibilities are depicted for natural language (top) and simple arithmetic (bottom). In either case, a stimulus-response learner (left) would simply memorize a lookup table where entries corresponded to entire inputs/outputs, ignoring any compositional structure. A compositional learner (right) would solve the same problems by composing familiar elements according to known rules, allowing generalization. Figure reproduced from Russin et al., (2021).
@@
In a seminal study on compositionality in modern DNNs, Lake and Baroni, (2018) leveraged these principles to develop the SCAN task, a sequence-to-sequence task where the model takes instructions as inputs (e.g., “walk left”), and must return the corresponding sequence of actions (e.g., TURN_LEFT WALK). To test compositional generalization, the researchers created a systematic difference between the train and test sets: one of the four primitive verbs (“jump”) was left out of the training set except in its simplest form (“jump” → JUMP). This created a scenario like the one discussed above, where a model could in principle infer the underlying grammar from all of the complicated constructions containing the other primitive verbs (e.g., “walk twice,” “look twice,” etc.), and combine this with knowledge of the action associated with “jump” (JUMP) to generalize to the analogous constructions containing “jump” (e.g., “jump twice”).
@@
The results showed that even though standard recurrent neural networks (RNNs) performed nearly perfectly when the train and test sets were randomly split, these same networks performed quite poorly on this compositional generalization test (about 1.2% accuracy reported in the original study). This was true for all of the models they tested, and follow-up work showed that other standard architectures such as convolutional neural networks (Dessì and Baroni,, 2019) and transformers (Keysers et al.,, 2020) exhibit similar reductions in performance on the compositional test (though less extreme in the case of the convolutional neural network). Analogous experiments in humans using a modified version of the same task have shown that humans perform quite well (Lake et al.,, 2019; Lake and Baroni,, 2023). A number of subsequent machine learning studies followed up on this work, creating similar tests of compositional generalization and showing that these standard DNNs fail to generalize compositionally in domains such as semantic parsing (Kim and Linzen,, 2020), string manipulation with a probabilistic context-free grammar (Hupkes et al.,, 2020), mathematics (Saxton et al.,, 2019), natural-language translation (Dankers et al.,, 2022), and grounded language understanding (Ruis et al.,, 2020), among others.
@@
Another (perhaps compatible) way of understanding these results is to take a statistical learning theory perspective. Neural networks are universal function approximators (Cybenkot,, 1989; Hornik,, 1991) that assume that the train data and test data are sampled from the same distribution, i.e., that the train and test data will be independent and identically distributed (i.i.d.). Viewed through this lens, the generalization behaviors that are taken to be diagnostic of compositionality (Evans,, 1985; Fodor and Pylyshyn,, 1988) are out-of-domain (o.o.d.) generalizations, and require extrapolation, rather than interpolation (Marcus,, 1998, 2018). For an unbiased learner like a standard neural network, this kind of generalization should be unlikely or impossible.
@@
The extent to which a learner is expected to generalize beyond its training distribution is the extent to which it must make additional assumptions about the world. These assumptions, called ‘inductive biases’ in statistical learning parlance, can take many forms in neural networks, including architectural inductive biases or extra regularization terms in the loss function (e.g., weight decay, a regularization term that encourages a network’s weights to be small). Standard neural networks do not make very many of these assumptions and thus tend to struggle when even the most superficial distributional shifts occur between train and test (Lake et al.,, 2017). For example, standard neural networks used in deep reinforcement learning agents trained on Atari games fail to generalize when superficial features of the game are changed (Kansky et al.,, 2017). It is not surprising, then, that they also fail in compositional generalization settings where there can be extreme distributional shifts from train to test.